OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6) Place the symbols on the grid to complete the equation.

OpenStudy (anonymous):

Slope = 1 Turn into point slope : y - 3 = 1(x-3) slope-intercept : y = x so that means the y-intercept = 0 standard form : x - y = 0

OpenStudy (anonymous):

it says its ronq

OpenStudy (anonymous):

what does "it" say the answer is?